# Flash Inference: Near Linear Time Inference for Long Convolution Sequence Models and Beyond

### v1

@@ 5.1 Integrating Flash Inference in real world setting (Hyena Architecture)
Flash Inference significantly speeds-up the end-to-end inference by up to 1.6$\times$ and the convolution-based mixer component of the Hyena architecture by 50$\times$ compared to the baselines as shown in Figures‘2(a) and 2(b).
@@
Figure 2: Real world Hyena experiments: (a) End-to-end inference time breakdown shows Hybird provides 1.6$\times$ speed-up over optimized baselines (b) Cumulative mixer time of Hybrid scales 50$\times$ better (c) Hybrid shows low variance in per-token response time except at the tokens positions where large tiles are computed.
@@ 5.2 $\tau$ Implementations
Algorithm 2 assumes an implementation of primitive $\tau$ that accounts for the contribution of a range of inputs to a range of outputs of a convolution. We considered $7$ different implementations but, here, we only present results of the ones on the Pareto Frontier - that is, those optimal for at least some $(B,N,D,U)$ setting. There are $4$ such implementations, of two types:
@@
    Depthwise-separable 1D Convolution: We use two types of implementations (a) Conv1D refers to the default PyTorch (Paszke et al., 2019) kernel, that requires explicit padding (b) Flash Conv1D refers to fused kernel by FlashFFTConv (Fu et al., 2023b) - these are asymptotically quadratic in tile size $U$.
@@
    FFT based convolution: we again use two implementations (a) PyTorch native, shown as FFT, that requires computing the FFTs of inputs, followed by pointwise multiplication and lastly performing the inverse FFT operation. (b) the second is provided by FlashFFTConv that does all the above in a single fused kernel shown as FlashFFT- both these scale as $O(U\log U)$.
@@ 5.3 Mixer Isolation study and Hybridization of $\tau$ implementations
We evaluate the different convolution implementations for all settings of $B,D,M,U$ and observe that each of these four implementation lies on the pareto frontier curve of tile size vs latency as shown in Figure 3(a). Our Hybrid approach dynamically chooses the best $\tau$ implementation for a given tile size $U$ based on the isolated empirically-measured efficiency of each implementation as shown in Figures 3(b). Figure 3(c) shows the cumulative token time breakdown for the mixer and non-mixer components of the synthetic setup for all our $\tau$ implementations. We observe an increase in the non-mixer components due to significant reduction in GPU kernel time that exposes the CPU kernel dispatch overhead for the MLP blocks in the non-mixer components.
Figure 3: Mixer Isolation in a Synthetic setting: (a) Different implementations of $\tau$ are optimal for different tile sizes creating a pareto optimal curve for Hybrid to choose, (b) Cumulative mixer inference of Hybrid achieves the best of all $\tau$ Implementations (c) End-to-end cumulative token inference breakdown: methods based on our tiling expose CPU overhead due to significantly reduced FLOPs and DRAM accesses
@@ 5.4 Improvements justification
    The dynamic choice of best $\tau$ implementation for given tile $U$ hybrid outperforming any method using a fixed implementation (Figure 3(b)).
@@
    Our engineering contributions: first, the DFT for the convolutional kernel is pre-computed ahead of time for $\log_{2}(L)-1$ tile sizes. Second, Flash-FFT configurations are pre-initialized for these tile sizes to maximize hardware performance. Third, right padding is used instead of left padding to reduce computation time by half. Fourth, properties of circular convolution are exploited to halven FFT length. Finally, tile calculations are parallelized across layers to saturate memory bandwidth for small tile computations and optimize computation for large tile computations, resulting in improved performance (one can notice improvements of $10-20$% even in the Eager and Lazy implementations alone)
@@ 6 Conclusion and Further Work
We propose a framework for performing inference in certain autoregressive sequence models. Among such models, LCSMs such as Hyena, are noteworthy: there, our framework provides an $O(L\log^{2}L)$ inference algorithm which, when run empirically, yields up to ×$\times 1.6$ time-efficiency improvement. The framework exploits a causal, fractal tiling that helps save on data movement and share computation. Moreover, it allows for almost-complete across-layers parallelization of $\mathrm{mixer}$-related workload.
@@ 7 Acknowledgments
Appendix A Proofs
Lemma.
Let $1\leq l\leq r\leq l^{\prime}\leq r^{\prime}\leq L$ represent ranges of lengths $L_{1}=r-l+1$ and $L_{2}=r^{\prime}-l^{\prime}+1$ of $y$ and $z$, respectively. There exists an FFT-based algorithm running in $O(L_{1}+L_{2})$ space and $O((L_{1}+L_{2})\log(L_{1}+L_{2}))$ time complexity that, given access to $y_{[l,r]}$, computes the aggregated contributions of all $y_{[l,r]}$ to all of $z_{[l^{\prime},r^{\prime}]}$. That is, for each $l^{\prime}\leq t\leq r^{\prime}$, it computes:
for all $l^{\prime}\leq t\leq r^{\prime}$.
Proof.
Consider performing a convolution between $y_{[l,r]}$ and $\rho_{[l^{\prime}-r,r^{\prime}-l]}$. It holds that for each $j\in[l^{\prime},r^{\prime}]$ and $i\in[l,r]$, $j-i$ is in the range of $\rho$. On the other hand, truncating the output of the convolution appropriately, one can keep only the corresponding outputs for each $j\in[l^{\prime},r^{\prime}]$. The overall size of the FFT is then given by the sum of the lengths, which is at most $r-l+1+(r^{\prime}-l)-(l^{\prime}-r)+1=2(r-l+1)+(r^{\prime}-l^{\prime}+1)-1=O((r-l+1)+(r^{\prime}-l^{\prime}+1))$. Since an FFT of order $L$ runs in $O(L\log L)$ and takes up $O(L)$ memory, the conclusion follows. ∎
Theorem.
Under the assumptions P.1 and P.2, one can generate $L=2^{P}$ tokens autoregressively by performing, per layer, $L-1$ black-box calls to ${\mathcal{A}}$ as well as $L$ more calls to $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$. Concretely, there are $L/2=2^{P-1}$ calls to ${\mathcal{A}}$ of length $1$ each, $L/4=2^{P-2}$ calls of length $2$ each and so on up to $1$ call of length $L/2$. Hence, neglecting the $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$ part, the overall time complexity per mixer layer is:
$\text{FLOPs}=\sum_{q=0}^{P-1}{2^{P-1-q}{\mathcal{T}}(2^{q},2^{q})}$
Proof.
Figure 1 shows why the tiling used in Algorithm 2 covers every pair of contributions exactly once and in the correct order (remember, we only assumed $\mathrm{agg}$ to be associative, not necessarily commutative).
Only part remaining to be proved is the right counting of calls to ${\mathcal{A}}$ per layer: calls of length $2^{r}$ happen whenever $2^{r}$ divides $i$ but $2^{r+1}$ does not - that is, for $i\in\{2^{r},3\cdot 2^{r},5\cdot r^{2},\ldots(2^{P-r}-1)\cdot 2^{r}\}$, so $2^{P-r-1}$ calls. ∎
@@ Appendix B Extension to Data-Dependent Filters
The reason why data-dependent filters are harder to deal with is that $\mathrm{cont}(y,i,j)=y_{i}\cdot\rho_{j-i}$ depends on both $y_{i}$ and on what $\rho_{j-i}$ depends on. By only assuming causality, we can only access $\rho_{j-i}$ after $z_{j-i-1}$ has been computed. This stops Algorithm 2 from working out-of-the-box since when $i$ is a power of $2$, in order to account for the contribution of $a_{\ell-1,[1,i]}$ to $a_{\ell,[i+1,2i]}$ one will need access to $\rho_{\ell,[1,2i-1]}$ which is not guaranteed to be accessible (only $\rho_{\ell,[1,i]}$ can be assumed to be known at this stage).
@@
0:  first token $a_{1,0}$, $a_{\cdot,0}$, $\rho_{\cdot,0}$, $\mathrm{block}_{[1,M]}$ and $\mathrm{sampler}$
1:  Output: All activations $a_{0},\ldots,a_{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
2:  Initialize $a$ with zeros outside $a_{\cdot,0}$
3:  for $i\leftarrow 1$ to $L-1$ do
@@
6:        Compute $\rho_{\ell,i}$ as a causal function of data $a_{\ell-1,[1,i]}$ as per model specification
@@
8:        $a_{\ell,i}=\mathrm{block}_{\ell}(a_{\ell,i}+a_{\ell-1,i}*\rho_{\ell,0}+a_{\ell-1,0}\odot\rho_{\ell,i})$
@@
11:           # downgrade the power of $2$ by half if $i+1$ is already a power of $2$
@@
13:           $a_{\ell,[2U,4U-2]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=CONV(a_{\ell-1,[U,2U-1]},\rho_{\ell,[U,2U-1]})$
@@
15:           $a_{\ell,[i+1,i+2U-1]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=CONV(a_{\ell-1,[U,2U-1]},\rho_{\ell,[i-U+1,i]})$
16:           $a_{\ell,[i+1,i+2U-1]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=CONV(\rho_{\ell,[U,2U-1]},a_{\ell-1,[i-U+1,i]})$
@@ Appendix C Implementation Improvements
In Algorithm 2, the calls to ${\mathcal{A}}$ all have a shape of ${\mathcal{A}}(y,i-U+1,i,i+1,i+U)$ where $U$ is a power of $2$ that divides $i$. Following the proof of Lemma 1, we note that to implement ${\mathcal{A}}$, we need to perform the convolution of a segment of length $U$ of $y$ and a prefix of length $2U$ of $\rho$. However, following the convolution, of the $13U-$ 1 outputs, indexed $[0,3U-2]$, we are only interested in the middle $U$ of them, namely indices $[U,2U-1]$. The canonical way of performing convolutions via FFT involves padding by enough $0s$ and using an order large enough to store the whole output which rounded up to the closest power of $2$ would mean an order $4U$ FFT. However, using a $2U$ FFT call, which will perform a cyclical convolution, is enough for our purposes, since the values of interest are not affected by the cyclicity - that is, when folding outputs at $[2U,3U-2]$ onto $[0,U-2]$, we do not intersect $[U,2U-1]$. Furthermore, there are only $\log L$ different lengths of prefixes of $\rho$ involved in these convolution so one could, in fact, cache the DFTs for these as a precomputation step, thus dropping the number of DFT’s per convolution from $3$ to $2$, speeding up by a further ×$\times 1.5$ factor.

### v2

@@ 5.1 Integrating Flash Inference in real world setting (Hyena Architecture)
[…]
@@
[…]
@@ 5.2 $\tau$ Implementations
Algorithm 2 assumes an implementation of primitive $\tau$ that accounts for the contribution of a range of inputs to a range of outputs of a convolution. We considered $7$ different implementations but, here, we only present results of the ones on the Pareto Frontier - that is, those optimal for at least some $(B,N,D,U)$ setting (the others are covered in Appendix C.1). There are $4$ such implementations, of two types:
@@
    Depthwise-separable 1D Convolution: we use two types of implementations, both asymptotically quadratic in tile size $U$:
    1. (a)
      Conv1D refers to the default PyTorch (Paszke et al., 2019) kernel, that requires explicit padding
    2. (b)
      Flash Conv1D refers to fused kernel by FlashFFTConv (Fu et al., 2023b)
@@
    FFT based convolution: we again use two implementations, only that this time they scale as $O(U\log U)$:
    1. (a)
      PyTorch native, shown as FFT, that requires computing the DFTs of inputs, followed by pointwise multiplication and lastly performing the inverse DFT operation
    2. (b)
      the second is provided by FlashFFTConv that does all the above in a single fused kernel shown as FlashFFT
@@ 5.3 Mixer Isolation study and Hybridization of $\tau$ implementations
We evaluate the different convolution implementations for all settings of $B,D,M,U$ and observe that each of these four implementation lies on the pareto frontier curve of tile size vs latency as shown in Figure 3(a). Our Hybrid approach dynamically chooses the best $\tau$ implementation for a given tile size $U$ based on the isolated empirically-measured efficiency of each implementation as shown in Figures 3(a). Figure 3(c) shows the cumulative token time breakdown for the mixer and non-mixer components of the synthetic setup for all our $\tau$ implementations. We observe the non-mixer component staying largely the same across approaches - this is made possible through the employment of CUDA Graphs as further discussed in Appendix C.2.
Figure 3: Mixer Isolation in a Synthetic setting: (a) Different implementations of $\tau$ are optimal for different tile sizes creating a pareto optimal curve for Hybrid to choose, (b) Cumulative mixer inference of Hybrid achieves the best of all $\tau$ Implementations (c) End-to-end cumulative token inference breakdown
@@ 5.4 Improvements justification
    The dynamic choice of best $\tau$ implementation for given tile $U$ - hybrid outperforming any method using a fixed implementation (Figure 3(b)).
@@
    Our engineering contributions: first, the DFT for the convolutional kernel is pre-computed ahead of time for $\log_{2}(L)-1$ tile sizes. Second, Flash-FFT configurations are pre-initialized for these tile sizes to maximize hardware performance. Third, right padding is used instead of left padding to reduce computation time by half. Fourth, properties of circular convolution are exploited to halven FFT length. Finally, tile calculations are parallelized across layers to saturate memory bandwidth for small tile computations and optimize computation for large tile computations, resulting in improved performance (one can notice improvements of 102010-20% even in the Eager and Lazy implementations alone)
@@ 6 Conclusion and Further Work
We propose a framework for performing inference in certain autoregressive sequence models. Among such models, LCSMs are noteworthy: there, our framework provides an $O(L\log^{2}L)$ inference algorithm which, when run empirically, yields up to $7.8\times$ end-to-end time-efficiency improvement. The framework exploits a causal, fractal tiling that helps save on data movement and share computation. Moreover, it allows for almost-complete across-layers parallelization of $\mathrm{mixer}$-related workload.
@@ 7 Acknowledgments
Appendix A Contextual discussion
In this section, we extend our discussion on the contextual relevancy of LCSMs and our method. We focus on the SISO (single-input-single-output) setting unless otherwise specified and assume that such a primitive is used across channels of the embedding dimensions to form a mixer.
As discussed in Section 2.3, LCSMs directly compute the filter $\rho$ to be used to map an input sequence $y\in\mathbb{R}^{L}$ to an output sequence $z\in\mathbb{R}^{L}$ via $z_{t}\triangleq\sum_{i=1}^{t}{y_{i}\cdot\rho_{t-i}}$. We assume here that $\rho\in\mathbb{R}^{L}$ although it could be longer since no entry past $L$th position matters. While $\rho$ is often underparameterized, one could, in principle, directly learn its entries (and, indeed, [Fu et al., 2023a] successfuly do that by using some extra regularization). However, even in underparameterized settings, where $\rho=f(\theta)$ for some $\theta\in\Theta$ with $\dim(\Theta)\ll L$, what filters $\rho$ we end up dealing with is a function of data and $f$: here, $f$ implicitly bakes in a certain bias. While depending on bias (such as decaying pattern of $\rho$ as suggested by [Li et al., 2022]) one may assume $\rho$ to have certain exploitable structures, our method is broadly applicable regardless of such structural properties. That makes Flash Inference be bias and data-independent. For completeness, we mention in Section A.2 previous works exploiting such structure - that is, ways to perform fast inference assuming certain properties of $\rho$ hold.
Since, as mentioned in Section 2, there is a direct connection between long convolutions and LTI SSMs, we cover this more extensively in Section A.1.
A.1 State Space Models
While SSMs and LCSMs can fundamentally represent the exact same class of models, they differ through their parameterization. Works such as [Poli et al., 2023, Li et al., 2022, Romero et al., 2021b, a, Shi et al., 2023, Fu et al., 2023a] try to characterize the convolution implicitly. On the other hand, LTI SSMs [Gu et al., 2021a, 2022, Ma et al., 2022, 2024, Gu et al., 2021b, Gupta et al., 2022, Mehta et al., 2022] define a recurrent system that is equivalent to a convolution and provide an alternative perspective in doing so. We define discrete SISO SSMs below following the notation of [Gu et al., 2021a].
State space models (SSM) can be thought of as linear RNNs. In its simplest form, SSM’s $\mathrm{mixer}$ for a SISO operator mapping $y\in\mathbb{R}^{L}$ to $z\in\mathbb{R}^{L}$ is defined via an extra sequence of hidden states $u:\mathbb{R}^{L\times D^{\prime}}$ where $D^{\prime}$ is the SSM’s dimension. For a fixed input $y:\mathbb{R}^{L}$, one defines:
$u_{t}=A(t)u_{t-1}+B(t)y_{t}$
$z_{t}=C(t)^{\top}u_{t}$
where $A(t):\mathbb{R}^{D^{\prime}\times D^{\prime}},B(t),C(t):\mathbb{R}^{D^{\prime}}$. While models like Mamba [Gu and Dao, 2023] let $A,B,C$ be functions of $y$ and, thus, vary with $t$, linear-time invariant (LTI) models assume $A$, $B$ and $C$ to be fixed parameters. Thus, the equations rewrite to:
$u_{t}=Au_{t-1}+By_{t}$
$z_{t}=C^{\top}u_{t}$
Opening up the recurrence, we get that:
$z_{t}=\sum_{i=0}^{t}{(C^{\top}A^{t-i}B)\cdot y_{i}}$
which corresponds to a convolution with the filter $\rho$ implicitly defined via $\rho_{i}=C^{\top}A^{i}B$. If $A$ is further assumed to be diagonal, we call the corresponding SSM diagonal. Gupta et al. [2022] showed that one does not lose practical model quality by assuming $A$ to be diagonal. When one tried to perform inference on an LTI SSMs, one could use:
  • the recurrent view: keep around only the last state $u_{t-1}$ and update it in-place. This is why the memory of performing inference with an SSM scales with $D^{\prime}$ rather than $L$. The time complexity of performing one step is $O(D^{\prime 2})$ in the general case, but as mentioned, one can resort to diagonal SSMs which drop this complexity to $O(D^{\prime})$.
  • convolutional view: precompute the equivalent convolution $\rho$ and use Flash Inference.
A.2 Comparison to Structural Filter Exploitation
Following our discussion in Section A.1, LTI SSMs become a candidate of structural property that can be assumed on a filter $\rho\in\mathbb{R}^{L}$. While if one is to let $D^{\prime}$ grow as much as $L$, it is provably possible to represent any filter $\rho$, it is a structural assumption to assume that a given filter $\rho$ can be represented (or well-approximated) by an LTI SSM of a particular dimension $D^{\prime}$ (and this is what [Massaroli et al., 2024] do). However, if the filter has this structure then we can use the recurrent view and pay a time cost of $D^{\prime}$ per token generation (assuming a diagonal SSM) while having our activation space drop by a factor of $L/D^{\prime}$. Alternatively, we can store all the activations and use Flash Inference to get an average token step for $O(\log^{2}(L)$ time-budget. Hence, at least asymptotically, if $D^{\prime}\in O(\log^{2}(L))$, one would prefer the recurrent view (and thus using the $D^{\prime}$-dimensional SSM equivalent to $\rho$). Past this point, our method is faster but still uses memory that scales up with $L$ rather than $D^{\prime}$ so which one to choose is a function of our memory constraints.
Another noteworthy line of work follows dilated convolutional models and was popularized through [Shi et al., 2023], inspired in turn by [Van Den Oord et al., 2016]. Paine et al. [2016] show how such models allow for linear time inference, while also needing to store all activations. Thus, granted such structure, Flash Inference is inferior and [Paine et al., 2016]’s approach should be used.
Finally, it is worth emphasizing that Flash Inference can accommodate data-dependent filters, while online exploitation of structure is often prohibitively expensive (as in the case of distilling $\rho$ into an LTI SSM following [Massaroli et al., 2024]). Moreover, the class of data-dependent convolution-based models is larger than that of LTI SSMs simply because of the linear-time invariant assumption: one cannot know the whole $A,B,C$ when only half of $\rho$ is unlocked (similarly to how data-dependent SSMs, such as [Gu and Dao, 2023], can represent more than LCSMs can). And while there are no major data-dependent causal LCSMs yet, works such as [Arora et al., 2023, Karami and Ghodsi, 2024] show the potential of such a research direction.
A.3 On Flash Inference’s prefilling and autoregressive generation stages
We have touched briefly in Section 2.3.1 on how our method deals with prompts. In particular, if one is given a prompt of size $P$, FFT will be employed to fill in all the contributions of $y_{[1,P]}$ to $z_{[1,L]}$. The way this translates to the complete architecture is by computing all of $a^{[0,M]}_{[1,P]}$ as well as the contributions of these to the future LPL-P positions. One can do so at a fraction of the cost of the forward pass, by just caching the appropriate activations and then zeroing out positions in the range $[P+1,L]$ before proceeding to the next layer (to not allow second order contributions to be performed). This comes at a cost that scales with $O(L\log L)$.
Afterwards, we can drop the first $P$ positions and apply our method starting with the current values of $a$ to perform autoregressive generation. This is because the eager nature of our prefill is more aggressive than in the case of transformers since we already account for all the relevant contributions of the prompt to the future. This can be very useful when we have a long prompt, relatively few autoregressive steps to perform and the autoregressive generation takes place on a different machine, since we save up valuable and communication costs.
We also want to emphasize, that Flash Inference focuses on optimizing the autoregressive generation part (also known as decoding). However, if the prompt is large enough compared to the number of tokens generated, this may have smaller relative improvement since the overall time is dominated by the prefill.
@@ Appendix B Extension to Data-Dependent Filters
The reason why data-dependent filters are harder to deal with is that $\mathrm{cont}(y,i,j)=y_{i}\cdot\rho_{j-i}$ depends on both $y_{i}$ and on what $\rho_{j-i}$ depends on. By only assuming causality, we can only access $\rho_{j-i}$ after $z_{j-i-1}$ has been computed. This stops Algorithm 2 from working out-of-the-box since when $i$ is a power of $2$, in order to account for the contribution of $a^{\ell-1}_{[1,i]}$ to $a^{\ell}_{[i+1,2i]}$ one will need access to $\rho^{\ell}_{[1,2i-1]}$ which is not guaranteed to be accessible (only $\rho^{\ell}_{[1,i]}$ can be assumed to be known at this stage).
@@
0: first token $a^{1}_{0}$, $a^{:}_{0}$, $\rho^{:}_{0}$, $\mathrm{block}^{[1,M]}$ and $\mathrm{sampler}$
1:Output: All activations $a^{0},\ldots,a^{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
2: Initialize $a$ with zeros outside $a^{:}_{0}$
3:for $i\leftarrow 1$ to L1L-1 do
@@
6:   Compute $\rho^{\ell}_{i}$ as a causal function of data $a^{\ell-1}_{[1,i]}$ as per model specification
@@
8:   $a^{\ell}_{i}=\mathrm{block}^{\ell}(a^{\ell}_{i}+a^{\ell-1}_{i}*\rho^{\ell}_{0}+a^{\ell-1}_{0}\odot\rho^{\ell}_{i})$
@@
11:    # downgrade the power of 22 by half if $i+1$ is already a power of 22
@@
13:    $a^{\ell}_{[2U,4U-2]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=CONV(a^{\ell-1}_{[U,2U-1]},\rho^{\ell}_{[U,2U-1]})$
@@
15:    $a^{\ell}_{[i+1,i+2U-1]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=CONV(a^{\ell-1}_{[U,2U-1]},\rho^{\ell}_{[i-U+1,i]})$
16:    $a^{\ell}_{[i+1,i+2U-1]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=CONV(\rho^{\ell}_{[U,2U-1]},a^{\ell-1}_{[i-U+1,i]})$
@@ Appendix C Implementation Improvements
In Algorithm 2, the calls to $\tau$ all have a shape of $\tau(y,[i-U+1,i],\rho_{,}[i+1,i+U])$ where $U$ is a power of $2$ that divides $i$. Following the proof of Lemma 1, we note that to implement $\tau$, we need to perform the convolution of a segment of length $U$ of $y$ and a prefix of length $2U$ of $\rho$. However, following the convolution, of the 3U13U-1 outputs, indexed [0,3U2][0,3U-2], we are only interested in the middle $U$ of them, namely indices [U,2U1][U,2U-1]. The canonical way of performing convolutions via FFT involves padding by enough $0s$ and using an order large enough to store the whole output which rounded up to the closest power of $2$ would mean an order $4U$ FFT. However, using a $2U$ FFT call, which will perform a cyclical convolution, is enough for our purposes, since the values of interest are not affected by the cyclicity - that is, when folding outputs at [2U,3U2][2U,3U-2] onto [0,U2][0,U-2], we do not intersect [U,2U1][U,2U-1]. Furthermore, there are only $\log L$ different lengths of prefixes of $\rho$ involved in these convolution so one could, in fact, cache the DFTs for these as a precomputation step, thus dropping the number of DFT’s per convolution from $3$ to $2$, speeding up by a further $\times 1.5$ factor.
C.1 More Implementation Details
In Section 5.2, it has been mentioned that we considered $7$ implementations of $\tau$. Besides the four mentioned (Torch’s and FlashFFT’s Conv1D and FFT) we also did 3 more naive versions:
  1. 1.
    Torch FFT (no precomputed FFT for the convolution kernel): This implementation is identical to Torch-FFT except that we do not pre-compute the DFT of the convolution kernel beforehand.
  2. 2.
    Vectorized Conv1D: It implements a 1D convolution operation in a vectorized manner using tensors and leverages efficient operations to process multiple slices in parallel. It takes advantage of the fact that convolutions are depth-wise separable.
  3. 3.
    Element-wise Conv1D: This implementation does an element-wise dot product of the convolution kernel with the input, while iteratively sliding the kernel over the input.
C.2 CPU Overhead: The silent performance killer
One primary improvement we introduce is the use of CUDA Graphs to mitigate the CPU overhead observed in the non-mixer components during inference.
In both synthetic settings and the Hyena model, these non-mixer components primarily deal with only current token being generated. As a result, the GPU compute time is often shorter than the kernel dispatch time, leading to CPU overhead. Typically, this overhead remains hidden when the mixer component requires significant GPU compute time. However, Flash Inference dramatically reduces mixer times (up to $120\times$), exposing the CPU overhead and causing the observed decrease in end-to-end speed-ups. A similar effect is observed for the Lazy approach compared to Eager, as Lazy’s faster execution also exposes the overhead.
To address this, we employ CUDA Graphs, which record all kernel dispatches for generating a single token into a graph. For subsequent token generation, the same graph is replayed, allowing all kernels for the non-mixer components to be dispatched simultaneously. This approach significantly reduces CPU overhead and ensures consistent overhead times for both Flash Inference and its baselines.
With these improvements, Flash Inference demonstrates much better end-to-end relative gains, as the mixer now constitutes a larger portion of the total runtime. We observe speed-ups of up to $120\times$ in the mixer and up to $8\times$ in the end-to-end Hyena runs (as opposed to $1.6\times$ when not using CUDA Graphs). All the results are shown in Appendix H.